OpenStudy (anonymous):

Can someone, with very detailed steps, show me how to answer the following 2x+1 over x+2 ; + 2x over x = 2x-3 over x-1

OpenStudy (anonymous):

[(2x+1)/(x+2)]+[2x/2]=[(2x-3)/(x-1)] is your equation look like this?

OpenStudy (anonymous):

@Bryan1707 ??? are you still there?

OpenStudy (anonymous):

Why's your middle term 2x/2 meyjoux. It says 2x/x.

OpenStudy (anonymous):

oops sorry .. mistakes

OpenStudy (anonymous):

[(2x+1)/(x+2)]+[2x/x]=[(2x-3)/(x-1)] first get the GCF of the following fraction. the GCF is( x+2)(x)(x-1) divide the GCF to all of the denominators and multiply the answer to all the numerators.. it will give you (2)(x+1)(2x+1)+(2x)(x+2)(x+1)=(2x-3)(x)(x+2) then perform multiplication to all the terms

OpenStudy (anonymous):

then simplify it..

OpenStudy (anonymous):

can you do that @Bryan1707 ?